# PreAlign: Boosting Cross-Lingual Transfer by Early Establishment of Multilingual Alignment

### v1

@@ Abstract
Large language models demonstrate reasonable multilingual abilities, despite predominantly English-centric pretraining. However, the spontaneous multilingual alignment in these models is shown to be weak, leading to unsatisfactory cross-lingual transfer and knowledge sharing. Previous works attempt to address this issue by explicitly injecting multilingual alignment information during or after pretraining. Thus for the early stage in pretraining, the alignment is weak for sharing information or knowledge across languages. In this paper, we propose PreAlign, a framework that establishes multilingual alignment prior to language model pretraining. PreAlign injects multilingual alignment by initializing the model to generate similar representations of aligned words and preserves this alignment using a code-switching strategy during pretraining. Extensive experiments in a synthetic English to English-Clone setting demonstrate that PreAlign significantly outperforms standard multilingual joint training in language modeling, zero-shot cross-lingual transfer, and cross-lingual knowledge application. Further experiments in real-world scenarios further validate PreAlign’s effectiveness across various model sizes.  [^1] The code of this paper is availabl at https://github.com/Saltychtao/PreAlign
@@
Jiahuan Li, Shujian Huang, Xinyu Dai  and Jiajun Chen National Key Laboratory for Novel Software Technology, Nanjing University, China lijh@smail.nju.edu.cn, {huangsj,daixinyu, chenjj}@nju.edu.cn
@@ 1 Introduction
A handful of works (Reimers and Gurevych, 2020; Cao et al., 2020; Wu and Dredze, 2020; Chaudhary et al., 2020; Yang et al., 2021; Tang et al., 2022; Feng et al., 2022; Gao et al., 2024) try to mitigate the problem by explicitly injecting alignment information using existing supervision data. They either construct cross-lingual prediction tasks (Chaudhary et al., 2020; Yang et al., 2021) or train models to produce similar representations of aligned words or sentences (Tang et al., 2022; Wu and Dredze, 2020; Reimers and Gurevych, 2020). Although these methods can bring reasonable improvements, the establishment of multilingual alignment requires a long training process either during or after pretraining (Dufter and Schütze, 2020), which prevents the model from effectively performing cross-lingual transfer at earlier stage in pretraining.
@@
More specifically, before large-scale language pretraining, PreAlign first collects multilingual translation pairs between English and languages to be transferred, and inject this information into the model by initializing it to produce similar representations of aligned words. In order to maintain the established multilingual alignment across the pretraining phase, we propose an input-only codeswitching strategy, which only substitutes words in the input text to its aligned words, and optimizes model using language modeling objective. The illustration of PreAlign is shown in Figure 1.
We firstly conduct experiments on a English to English-Clone settings (K et al., 2020; Dufter and Schütze, 2020; Schäfer et al., 2024). English-clone is a synthetic language that shares identical grammar and vocabulary distribution with English, but no vocab overlap. This allows us to study cross-lingual transfer on a more controlled environment. Experiments demonstrate that PreAlign significantly improves models’ ability of languages to be transferred, and strengthens cross-lingual transfer of downstream task abilities and knowledge. Further analysis shows that the early established multilingual alignment can be kept throughout large-scale language pretraining and generalize to other words. We further experiment with our methods on real-world settings, and validates the effectiveness of PreAlign across different model scales.
@@ 2.1 Understanding Cross-lingual Ability of Pretrained language models
Many works attempt to analyze the cross-lingual ability of LLMs. Dufter and Schütze (2020); Conneau et al. (2020) try to explain factors that contributes to spontaneous multilingual alignment developed in pretrained language models, including under-parameterization, shared model architectures and pivot words across languages. Other works investigate the working mechanism of multilingual representations. Wendler et al. (2024) find that English-centric models works on a concept space that is close to English when processing other languages. Recently, Gao et al. (2024); Qi et al. (2023) analyze multilingual knowledge alignment in existing LLMs, and find that multilingual training and instruction tuning can only lead to shallow alignment, i.e. LLMs can achieve similar task performances and consistent responses across languages, yet cannot apply knowledge across languages.
@@ 3.1 Injecting Multilingual Alignment before Language Pretraining
Given an English monolingual corpus $\mathcal{D}$, PreAlign extracts from $\mathcal{D}$ the collections of all unique words $\mathcal{W}=\{w\}_{i}^{N}$, where $N$ is the number of unique words. For each word $w$, we translate it to all considered target languages, and denote the translation results as $T(w)$. Since there exist complex many-to-many alignment relationships between languages, PreAlign needs to collect all possible translations. We rely on GPT-4 to collect the corresponding translations in this paper.
@@
After the multilingual alignment table is collected, PreAlign initializes models’ parameters using a contrastive alignment objective, which optimizes the model to produce similar representations for aligned words. Specifically, given an English word $w_{i}$ and its translations across all other languages $T(w_{i})$, PreAlign firstly obtains representations of each layer for each $w\in\mathcal{S}_{w_{i}}$:
@@
where $l=0,1,\cdots,L,L+1$. $f(w,l),1\leq l\leq L$ denotes of the $l$-th Transformer layer representations of the model’s encoding of $w$. $f(w,0)$ and $f(w,L+1)$ denotes the word embedding and output embedding of $w$, respectively. Note that since $w$ could be tokenized to multiple subwords, PreAlign aggregates them into a single representation using mean-pooling operator.
@@
$\mathcal{L}_{align}=\sum_{l=0}^{L+1}\mathcal{L}_{align}^{l}$ (3)
To prevent the initialization from being trapped in a local minima that is not suitable for the subsequent language modeling, we also add an auxiliary language modeling loss beside the contrastive objective in practice:
$\mathcal{L}_{joint}=\mathcal{L}_{align}+\mathcal{L}_{LM}$ (4)
@@ 3.2 Maintaining Multilingual Alignment via Input-only Codeswitching
The method described previously introduces multilingual alignment information before language pretraining. However, this information may be quickly forgotten if not continuously reinforced. Inspired by prior research  (Chaudhary et al., 2020; Yang et al., 2021) demonstrating that code-switching effectively promotes multilingual alignment, we propose using the code-switching technique to sustain this alignment throughout the pretraining process.
Originally, code-switching was applied to both the input sequence and the target tokens in raw data, posing no issues for pretraining encoder-only models. However, this approach exacerbates the issue of multilingual script mixing in the outputs of decoder-only models. To address this, we propose an input-only codeswitching strategy that affects only the input. The distinction between the traditional codeswitching and our input-only codeswitching is illustrated in Figure 2.
Formally, given a subword sequence $X_{<i}x_{i}^{1}\cdots x^{i}_{m}X_{>i}$, where $X_{<i}$ and $X_{>i}$ are the subword sequences before and after the $i$-th word, respectively. $x_{i}^{1}\cdots x_{i}^{m}$ is the subword sequence of the $i$-th words. Suppose the $i$-th word is substituted by $y_{i}^{1}\cdots y_{i}^{n}$ after codeswitching, then the language modeling objective after the original codeswitching is
$p(X_{<i})$ $\cdot p(X_{>i}|y_{i}^{1}\cdots y_{i}^{n})$
$\cdot p(y_{i}^{1}|X_{<i})\cdot\prod_{j=2}^{n}p(y_{i}^{j}|X_{<i}y_{i}^{1}\cdots y_{i}^{j-1})$ (5)
In Equation 5, the item $p(y_{i}^{1}|X_{<i})$ requires the model to generate words in another language given prefixes in one language. To mitigate this, input-only codeswitching modifies the objective to be
$p(X_{<i})\cdot p(X_{>i}|y_{i}^{1}\cdots y_{i}^{n})\cdot p(x_{i}^{1}|X_{<i}).$ (6)
Equation 6 changes the prediction objective of subwords in the word after codeswitching ($p(y_{i}^{1}|X_{<i})$) to subwords in the word before codeswitching ($p(x_{i}^{1}|X_{<i})$), therefore preventing the generation results contain scripts from other languages. In this paper, we use a codeswitching ratio of 5%.
@@
4 Experimental Settings
4.1 Datasets and Models
@@
We adopt the GPT-2 style Transformer architecture for our model. At the defaulting setting, our model contains 12 Transformer layers with a hidden dimension of 1024. The number of total non-embedding parameters is about 150 million. We use AdamW (Kingma and Ba, 2017) optimizer with a global batch size of about 1 million tokens. The learning rate is decayed from $43e-$ 4 to $53e-$ 5 following a cosine scheduler.
Pretraining Dataset
We adopt CulturaX (Nguyen et al., 2023) as the pretraining dataset. CulturaX is a multilingual pretraining corpus that has been rigorously cleaned. Due to the non-affordable computational cost to use all data for experiments, we only consider English as the source language, and Chinese (Zh), German (De), Russian (Ru), Arabia (Ar) as the target language. For English, we randomly select 10 billion tokens from CulturaX as the pretraining data. For each language to be transferred to, we randomly select 100 million tokens.
@@
Table 1: Performance of PreAlign and other methods on language modeling, ZS-CLT and CLKA. The performance of CLKA is averaged over each learning period.
4.2 Evaluation Protocol
Target Language Modeling (LM)
The first evaluation metric is the language modeling performance of target language. Given the same amount of target language data, this can reflect how well cross-lingual transfer is.
Zero-shot Cross-lingual Transfer (ZS-CLT)
Another common way to evaluate model’s cross-lingual ability is zero-shot cross-lingual transfer, where we finetune models on the task data in source languages, and test model’s ability on the same task in target languages. We use the commonly-used XNLI (Conneau et al., 2018) dataset for ZS-CLT evaluation.
Cross-lingual Knowledge Application (CLKA)
Large language models acquire extensive world knowledge from their pretraining corpora. However, significant portions of knowledge exist exclusively in texts of specific languages. It is ideal for LLMs to learn knowledge from texts in one language and apply it across other languages.
In order to evaluate models’ ability to perform such cross-lingual knowledge application, we propose a setting where we attach English texts describing synthetic knowledge to the pretraining corpus, and test models’ completion accuracy of the injected knowledge in the target language. Each synthetic knowledge is a triplet like (subject, relation, object), where relations are extracted from WikiData (Vrandečić and Krötzsch, 2014), and subjects and objects are artificial entities.
To better monitor the model’s learning dynamics, we segmented the pretraining process into shorter periods, each consisting of 250 training steps. During each period, we incorporate various knowledge triplets into predefined templates to create sentences that encapsulate specific knowledge, which are then added to the pretraining data exclusively during that period. Following each learning period, we assess the model’s knowledge retention by introducing three distractors—random named entities substituted for the original object in the knowledge statement—and evaluate the model’s ability to correctly assign the highest likelihood to the correct statement. This assessment occurs immediately after each training period using the corresponding model checkpoint.
5 Experiments on Synthetic Transferring Settings
Figure 3: Illustration of the creation of English-Clone.
@@ 5 Experiments on Synthetic Transferring Settings
We start our evaluation on a English to Synthetic language transferring setting, which allows us to better control the relationship between the source language and target language. We first describe the construction of synthetic language and implication of the setting in Section 5.1. We then present experimental results in Section 5.2 and Section 5.3. Finally, in-depth discussions are presented in Section 5.4,  5.5 and  5.6.

### v2

@@ Abstract
Large language models demonstrate reasonable multilingual abilities, despite predominantly English-centric pretraining. However, the spontaneous multilingual alignment in these models is shown to be weak, leading to unsatisfactory cross-lingual transfer and knowledge sharing. Previous works attempt to address this issue by explicitly injecting multilingual alignment information during or after pretraining. Thus for the early stage in pretraining, the alignment is weak for sharing information or knowledge across languages. In this paper, we propose PreAlign, a framework that establishes multilingual alignment prior to language model pretraining. PreAlign injects multilingual alignment by initializing the model to generate similar representations of aligned words and preserves this alignment using a code-switching strategy during pretraining. Extensive experiments in a synthetic English to English-Clone setting demonstrate that PreAlign significantly outperforms standard multilingual joint training in language modeling, zero-shot cross-lingual transfer, and cross-lingual knowledge application. Further experiments in real-world scenarios further validate PreAlign’s effectiveness across various languages and model sizes.  [^1] The code of this paper is available at https://github.com/Saltychtao/PreAlign
@@
Jiahuan Li, Shujian Huang†♣, Aarron Ching, Xinyu Daiand Jiajun Chen National Key Laboratory for Novel Software Technology, Nanjing University, China Independent Researcher lijh@smail.nju.edu.cn, {huangsj,daixinyu, chenjj}@nju.edu.cn
@@ 1 Introduction
A handful of works (Reimers and Gurevych, 2020; Cao et al., 2020; Wu and Dredze, 2020; Chaudhary et al., 2020; Yang et al., 2021; Tang et al., 2022; Feng et al., 2022; Gao et al., 2024) try to mitigate the problem by explicitly injecting alignment information using existing supervision data. They either construct cross-lingual prediction tasks (Chaudhary et al., 2020; Yang et al., 2021) or train models to produce similar representations of aligned words or sentences (Tang et al., 2022; Wu and Dredze, 2020; Reimers and Gurevych, 2020). However, the improvements are somewhat mixed and the establishment of multilingual alignment requires a long training process either during or after pretraining (Dufter and Schütze, 2020), which prevents the model from effectively performing cross-lingual transfer at earlier stage in pretraining.
@@
More specifically, before large-scale language pretraining, PreAlign first collects multilingual translation pairs between English and languages to be transferred, and inject this information into the model by pre-pretraining it to produce similar representations of aligned pairs. In order to maintain the established multilingual alignment across the pretraining phase, we propose an input-only codeswitching strategy, which only substitutes words in the input text to its aligned words, and optimizes model using language modeling objective. The illustration of PreAlign is shown in Figure 1.
We firstly conduct experiments on an English to English-Clone setting (K et al., 2020; Dufter and Schütze, 2020; Schäfer et al., 2024), where English-clone is a synthetic language that shares identical grammar and vocabulary distribution with English, but has no vocab overlap. This allows us to study cross-lingual transfer on a more controlled environment. Experiments demonstrate that PreAlign improves language ability on English-Clone by strengthening the cross-lingual transfer of knowledge and abilities from English. Further analysis shows that the early established multilingual alignment can be kept throughout large-scale language pretraining and generalize to other unaligned words. Experiments on real-world settings (including Chinese, German, Arabic and Russian) validate the effectiveness of PreAlign across different languages and model scales.
@@ 2.1 Understanding Cross-lingual Ability of Pretrained language models
Many works attempt to analyze the cross-lingual ability of LLMs. Dufter and Schütze (2020); Conneau et al. (2020) try to explain factors that contributes to spontaneous multilingual alignment developed in pretrained language models, including under-parameterization, shared model architectures and pivot words across languages. Other works investigate the working mechanism of multilingual representations. Wendler et al. (2024) find that English-centric models works on a concept space that is close to English when processing other languages. Gaschi et al. (2023); Hämmerl et al. (2024) discuss the relationship between multilingual alignment and cross-lingual transfer. Recently, Gao et al. (2024); Qi et al. (2023) analyze multilingual knowledge alignment in existing LLMs, and find that multilingual training and instruction tuning can only lead to shallow alignment, i.e. LLMs can achieve similar task performances and consistent responses across languages, yet cannot apply knowledge across languages.
@@ 3.1 Injecting Multilingual Alignment before Language Pretraining
The collection of multilingual alignment table can either leverage the off-the-shelf multilingual dictionaries such as MUSE, or rely on machine translation models. In this paper, we take the second method: we first extract from an English monolingual corpus $\mathcal{D}$ the collections of all unique words $\mathcal{W}=\{w\}_{i}^{N}$, where $N$ is the number of unique words. For each word $w$, we translate it to all considered target languages, and denote the translation results as $T(w)$. We collect diverse translations for each word using GPT-4. More details can be found in Appendix A.
@@
After the multilingual alignment table is collected, PreAlign initializes models’ parameters using a contrastive alignment objective, which optimizes the model to produce similar representations for aligned words. Specifically, given an English word $w_{i}$ and its available translations across other languages $T(w_{i})$, PreAlign firstly obtains representations of each layer for each $w\in T(w_{i})$:
@@
where $l=0,1,\cdots,L,L+1$; $f(w,l)$ for $1\leq l\leq L$ denotes the $l$-th Transformer layer representations of the model’s encoding of $w$; $f(w,0)$ and $f(w,L+1)$ denotes the word embedding and output embedding of $w$, respectively. Note that since $w$ could be tokenized to multiple subwords, PreAlign aggregates them into a single representation using mean-pooling operator.
@@
$\mathcal{L}_{\text{align}}=\sum_{l=0}^{L+1}\mathcal{L}_{\text{align}}^{l}$ (3)
To prevent the initialization from being trapped in a local minima that is not suitable for the subsequent language modeling, we also add an auxiliary language modeling loss beside the contrastive objective in practice  [^2] The training data for $\mathcal{L}_{\text{align}}$ and $\mathcal{L}_{\text{LM}}$ are independently sampled in each mini-batch.:
$\mathcal{L}_{\text{joint}}=\mathcal{L}_{\text{align}}+\mathcal{L}_{\text{LM}}$ (4)
@@ 3.2 Maintaining Multilingual Alignment via Input-only Codeswitching
PreAlign injects multilingual alignment information before language pretraining. However, it is possible that this information could be quickly forgotten if not continuously reinforced. Inspired by prior research  (Chaudhary et al., 2020; Yang et al., 2021) demonstrating that codeswitching effectively promotes multilingual alignment, we propose using the codeswitching technique to sustain this alignment throughout the pretraining process.
Originally, codeswitching was applied to both the input sequence and the target tokens in raw data, which can exacerbate the issue of multilingual script mixing in the outputs of decoder-only models. To address this, we propose an input-only codeswitching strategy that affects only the input. The distinction between the traditional codeswitching and our input-only codeswitching is illustrated in Figure 2.
Formally, given a subword sequence $X_{<i},x_{i}^{1},\cdots,x^{i}_{m},X_{>i}$, where $X_{<i}$ and $X_{>i}$ are the subword sequences before and after the $i$-th word, respectively; $x_{i}^{1},\cdots,x_{i}^{m}$ is the subword sequence of the $i$-th words. Suppose the $i$-th word is substituted by $y_{i}^{1},\cdots,y_{i}^{n}$ after codeswitching, then the language modeling objective after the original codeswitching is
$\mathcal{L}_{\text{vanilla\_CS}}=p(X_{<i})$ $\cdot p(X_{>i}|X_{<i},y_{i}^{1},\cdots,y_{i}^{n})$
$\cdot p(y_{i}^{1}|X_{<i})$
$\cdot\prod_{j=2}^{n}p(y_{i}^{j}|X_{<i},y_{i}^{1},\cdots,y_{i}^{j-1})$ (5)
In Equation 5, the item $p(y_{i}^{1}|X_{<i})$ requires the model to generate words in one language given prefixes in another language. To mitigate this, our input-only codeswitching modifies the objective to be
$\mathcal{L}_{\text{input\_only\_CS}}=p(X_{<i})$ $\cdot p(X_{>i}|X_{<i},y_{i}^{1},\cdots,y_{i}^{n})$
$\cdot p(x_{i}^{1}|X_{<i}).$ (6)
Equation 6 omits the prediction objective of subwords in the word after codeswitching ($p(y_{i}^{1}|X_{<i})$), therefore preventing the generation results contain scripts from other languages. In this paper, we use a codeswitching ratio of 5%.
@@
4 Evaluation of Cross-Lingual Transfer
To investigate the cross-lingual transfer effects that is close to situations in current LLMs, we design the evaluation in an English-dominated setting, where most of the pretraining data is English. For the examined target language, the amount of pretraining data is much less than English. Intuitively, the language ability of the target language will be much weaker than English. However, it is still interesting to know to what extent the language abilities and knowledge could transfer from English to the target language.
4.1 Languages
Figure 3: Illustration of the creation of English-Clone.
4.1.1 Synthetic Language: En-Clone
We construct a synthetic language called En-Clone, by cloning all English tokens by a one-to-one mapping. En-Clone shares the same linguistic properties with English, such as vocabulary distribution, subword segmentation, grammar and syntax, yet they have no word overlapping. See Figure 3 for an illustration.
4.1.2 Real-World Languages
We also experiment with real-world languages to examine the effect of PreAlign in more complex situations. We select Chinese, Russian, German and Arabic as our target languages, which spans four different language families, serving as good representatives of world languages. Note that in this case, the alignment is established for 4 target languages at the same time.
4.2 Evaluation Metrics
The cross-lingual transfer effects are evaluated in the following 3 aspects:
Target Language Modeling (LM)
The first evaluation metric is the language modeling performance (perplexity) of the target language. Given the same amount of target language data, this can reflect the extent of language ability transferred from English to the target language.
Zero-shot Cross-lingual Transfer (ZS-CLT)
Another common way to evaluate model’s cross-lingual ability is zero-shot cross-lingual transfer, where we finetune models with training data of a given task in the source language, and test model’s ability on the same task in target languages. We use the commonly-used XNLI (Conneau et al., 2018) dataset for ZS-CLT evaluation.
Cross-lingual Knowledge Application (CLKA)
Large language models acquire extensive world knowledge from their pretraining corpora, which might be described in different languages. It is ideal for LLMs to learn knowledge from texts in one language and apply it across other languages.
In order to evaluate models’ ability to perform such cross-lingual knowledge application, we propose a setting where the model is trained with certain English texts describing synthetic knowledge, and test the injected knowledge in the target language. Each synthetic knowledge is a triplet like (subject, relation, object), where relations are extracted from WikiData (Vrandečić and Krötzsch, 2014), and subjects and objects are artificial entities. We assess the model’s knowledge retention by comparing the likelihood of different statements, including one correct statement and three distractors generated by randomly substituting named-entities for the original object in the knowledge statement. See Appendix A for examples of synthesized knowledge.
4.3 Experiment Settings in General
Pretraining Dataset
We adopt CulturaX (Nguyen et al., 2023) as the pretraining dataset. CulturaX is a multilingual pretraining corpus that has been rigorously cleaned. For English, we randomly select 10 billion tokens from CulturaX as the pretraining data. For each language to be transferred to (Zh, De, Ru and Ar in the real-world setting, and En-Clone in the synthetic setting), we randomly select 100 million tokens, which is 1% of the data in English.
@@ 4.3 Experiment Settings in General
We adopt the GPT-2 style Transformer architecture for our model. As the defaulting setting, our model contains 12 Transformer layers with a hidden dimension of 1024. The number of total non-embedding parameters is about 150 million. We use AdamW (Kingma and Ba, 2017) optimizer with a global batch size of about 1 million tokens. The learning rate is decayed from $43e-$ 4 to $53e-$ 5 following a cosine scheduler.
Baselines
We compare PreAlign ’s performance with the following methods:
  • Joint Training, where we pretrain the model on the mix of 10 billion English tokens and 100 million tokens in the target language.
  • Only Target, where we only pretrain the model on 100 million tokens in the target language.
  • Full Target, where we pretrain the model on 10 billion tokens in the target language. This can serve as an upper-bound performance for the target language.
@@
Table 1: Performance of PreAlign and other methods on language modeling, zero-shot cross-lingual transfer (ZS-CLT) and cross-lingual knowledge application (CLKA).
5 Experiments on Synthetic Setting
@@
e start our evaluation on the English to En-Clone setting, which allows us to better control the relationship between the source and target language.
5.1 General Results
We present results on LM, ZS-CLT and CLKA in Table 1.
Table 1 shows that compared to Only-Target, Joint training achieves notable improvements on LM despite there are neither parallel signal or pivot words between English and English-clone. Surprisingly, the model could successfully transfer the ability to perform NLI task from English (79.8) to English-Clone (74.9). However, this transfer does not work well on CLKA, which is consistent with previous findings (Gao et al., 2024) that cross-lingual knowledge transfer is hard to achieve by multilingual pretraining.
PreAlign improves over Joint Training on all evaluation tasks.
We can also see that PreAlign significantly outperforms Joint Training on all three evaluation metrics. On the LM evaluation, PreAlign even achieves performance comparable to Full-Target, using only 1% data. For ZS-CLT, the performance gap between two languages are narrowed. For the CLKA the accuracy is greatly improved (from 27.7 to 64.6). All the results demonstrate the effectiveness of PreAlign for facilitating cross-lingual transfer. It is worth noticing that the performance of English are also improved, suggesting the learning of English-Clone also helps English as well.
PreAlign outperforms methods that establish alignment during and post pretraining.
We experiment with performing contrastive alignment during and post the pretraining process (Wu and Dredze, 2020), and compared them to PreAlign in Table 2. For the post-pretraining alignment, we add an additional LM loss to reduce catastrophic forgetting. It can be seen that on-the-fly alignment can degrade model’s performance, which we hypothesize that excessively optimizing the model on the limited size of word-level multilingual alignment during pretraining might have a negative impact on language ability. For post-pretraining alignment, we can observe a improvement over Joint Training on LM and ZS-CLT, but nearly no effect on CLKA. However, PreAlign outperforms both on-the-fly alignment and post alignment on all three evaluation protocols.
LM ZS-CLT CLKA
Joint Training 21.6 74.9 27.7
On-the-fly alignment 22.1 74.3 26.5
Post alignment 19.7 75.5 28.4
PreAlign 16.5 79.3 90.3
Table 2: Comparison of performing contrastive alignment at different stage. On-the-fly alignment: performing alignment during the pretraining. Post alignment: performing alignment when the pretraining is done.
5.2 In-Depth Investigation for CLKA
As the performance for CLKA varies for different methods, we further examine the learning dynamics of CLKA. We segment the pretraining process into shorter periods, each consisting of 250 training steps, and evaluate the CLKA accuracy for each period. More specifically, for each period, knowledge of different frequence are provided to the model, and assessment occurs immediately after each period using the corresponding model checkpoint. For comparison, we evaluate all four combinations of languages for training and testing the knowledge. Figure 4 shows the results.